Andrea and Brad live 18km apart. Brad walked towards Andrea's house at 1.5km/h. Four hours later, Andrea leaves home and walks towards Brad's house at 4.5km/h. Find when and where they meet.
Found 3 solutions by ankor@dixie-net.com, FahadXpro, MathTherapy:
Answer by ankor@dixie-net.com(22740)

Andrea and Brad live 18km apart.
Brad walked towards Andrea's house at 1.5km/h.
Four hours later, Andrea leaves home and walks towards Brad's house at 4.5km/h.
Find when and where they meet.
:
let t = B's walking time
A left four hrs later, therefore
(t-4) = A's walking time
;
When they meet, the total of their distances will be 18 km
write a distance equation; dist = speed * time
1.5t + 4.5(t-4) = 18
1.5t + 4.5t - 18 = 18
6t = 18 + 18
6t = 36
t = 36/6
t = 6 hrs is Brad's walking time when they meet
Andrea will have walked 4 hrs less or 2 hrs
:
Find the point from Brad's house where they meet
1.5(6) = 9 mi or half way
;
:
Check this, B walked 9 mi and A walked 4.5(2), 9 mi, a total of 18 mi

Answer by FahadXpro(1)

Distance between houses = 18km
Brad's speed = 1.5km/h
Andrea's speed = 4.5km/h
Delay in time = 4hrs
Time that brad will take if he went to Andreas house:
time = distance/speed
"""Put in the values"""
time = 18km/(1.5km/h)
"""Solve"""
time = 12hrs
Time Andrea will take if she went to Brads house:
time = distance/speed
"""Put in the values"""
time = 18km/(4.5km/h)
"""Solve"""
time = 4hrs
We know that Brad has traveled for 4hrs as Andrea started to walk 4hrs later so:
Distance brad traveled in 4hrs:
Distance = Speed*Time
"""Put in the values"""
Distance = 1.5km*4hrs
"""Solve"""
Distance = 6km
We know that they meet each other so Andrea will have traveled 18km - 6km = 12km so:
Time = Distance/speed
"""Put in the values"""
time = 12km/(4.5km/h)
"""Solve"""
time = (8/3)hrs
Brad walked 6km away from his home and Andrea walked 12km away from her house to meet each other
Brad took 4hrs
Andrea took (8/3)hrs
It can be wrong. You should confirm by checking the answers in the back of your book. I am just a beginner.

Answer by MathTherapy(10552)

Andrea and Brad live 18km apart. Brad walked towards Andrea's house at 1.5km/h. Four hours later, Andrea leaves home and walks towards Brad's house at 4.5km/h. Find when and where they meet.
When will they meet? 
Where will they meet?